# Finite Sample Analysis of Distributional TD Learning with Linear Function Approximation

### v1

@@ Appendix D Omitted Results and Proofs in Section 4
${\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}\preccurlyeq\frac{1}{(1-\sqrt{\gamma})^{2}}\bar{{\bm{A}}}^{\intercal}\left({{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}^{-1}}\right)\bar{{\bm{A}}}\left({\preccurlyeq\frac{4}{(1-\gamma)^{2}\lambda_{\min}}\bar{{\bm{A}}}^{\intercal}\bar{{\bm{A}}}}\right),$ (25)
@@
$\left({{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}}\right)\bar{{\bm{A}}}^{\intercal}\left({{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}^{-1}}\right)\bar{{\bm{A}}}\left({{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}}\right)\succcurlyeq\left({1-\sqrt{\gamma}}\right)^{2}.$
@@
$\bar{{\bm{A}}}$ $=\left({{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}\right)-{\mathbb{E}}_{s,r,s^{\prime}}\left[{\left({{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right)\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}}\right)}\right],$
then for any ${\bm{w}}\in{\mathbb{R}}^{dK}$ with $\left\|{{\bm{w}}}\right\|=1$,
${\bm{w}}^{\intercal}\left({{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}}\right)\bar{{\bm{A}}}^{\intercal}\left({{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}^{-1}}\right)\bar{{\bm{A}}}\left({{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}}\right){\bm{w}}$
$\qquad=\left\|{\left({{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}}\right)\bar{{\bm{A}}}\left({{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}}\right){\bm{w}}}\right\|^{2}$
$\qquad=\left\|{{\bm{w}}-{\mathbb{E}}_{s,r,s^{\prime}}\left[{\left({{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right)\otimes\left({{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}}\right)}\right]{\bm{w}}}\right\|^{2}$
$\qquad\geq\left({1-\left\|{{\mathbb{E}}_{s,r,s^{\prime}}\left[{\left({{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right)\otimes\left({{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}}\right)}\right]{\bm{w}}}\right\|}\right)^{2}.$
@@
$\left\|{{\mathbb{E}}_{s,r,s^{\prime}}\left[{\left({{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right)\otimes\left({{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}}\right)}\right]}\right\|\leq\sqrt{\gamma}.$ (26)
@@
$\left\|{{\mathbb{E}}_{s,r,s^{\prime}}\left[{{\bm{Y}}(r)\otimes\left({{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}}\right)}\right]}\right\|\leq\sqrt{\gamma}.$
For any vectors ${\bm{w}}=\left({{\bm{w}}(0)^{\intercal},\cdots,{\bm{w}}(K-1)^{\intercal}}\right)$ and ${\bm{v}}=\left({{\bm{v}}(0)^{\intercal},\cdots,{\bm{v}}(K-1)^{\intercal}}\right)$ in ${\mathbb{R}}^{dK}$, we define the corresponding matrices ${\bm{W}}=\left({{\bm{w}}(0),\cdots,{\bm{w}}(K-1)}\right)$ and ${\bm{V}}=\left({{\bm{v}}(0),\cdots,{\bm{v}}(K-1)}\right)$ in ${\mathbb{R}}^{d\times K}$, then $\left\|{{\bm{w}}}\right\|=\left\|{{\bm{W}}}\right\|_{F}=\sqrt{\operatorname{tr}\left({{\bm{W}}^{\intercal}{\bm{W}}}\right)}$ and $\left\|{{\bm{v}}}\right\|=\left\|{{\bm{V}}}\right\|_{F}=\sqrt{\operatorname{tr}\left({{\bm{V}}^{\intercal}{\bm{V}}}\right)}$. With these notations, we have
$\left\|{{\mathbb{E}}_{s,r,s^{\prime}}\left[{{\bm{Y}}(r)\otimes\left({{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}}\right)}\right]}\right\|$
$\qquad=\sup_{\left\|{{\bm{w}}}\right\|=\left\|{{\bm{v}}}\right\|=1}{\bm{w}}^{\intercal}{\mathbb{E}}_{s,r,s^{\prime}}\left[{{\bm{Y}}(r)\otimes\left({{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}}\right)}\right]{\bm{v}}$
$\qquad=\sup_{\left\|{{\bm{w}}}\right\|=\left\|{{\bm{v}}}\right\|=1}{\mathbb{E}}_{s,r,s^{\prime}}\left[{\sum_{i,j=1}^{K}y_{ij}(r){\bm{w}}(i)^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{v}}(j)}\right],$
@@
${\bm{w}}(i)^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{v}}(j)=\left({{\bm{W}}^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{V}}}\right)_{ij},$
@@
$\left\|{{\mathbb{E}}_{s,r,s^{\prime}}\left[{{\bm{Y}}(r)\otimes\left({{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}}\right)}\right]}\right\|$
$\qquad=\sup_{\left\|{{\bm{W}}}\right\|_{F}=\left\|{{\bm{V}}}\right\|_{F}=1}{\mathbb{E}}_{s,r,s^{\prime}}\left[{\sum_{i,j=1}^{K}y_{ij}(r)\left({{\bm{W}}^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{V}}}\right)_{ij}}\right]$
@@
$\qquad=\sup_{\left\|{{\bm{W}}}\right\|_{F}=\left\|{{\bm{V}}}\right\|_{F}=1}{\mathbb{E}}_{s,r,s^{\prime}}\left[{\operatorname{tr}\left({{\bm{Y}}(r){\bm{V}}^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s^{\prime}){\bm{\phi}}(s)^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{W}}}\right)}\right]$
$\qquad=\sup_{\left\|{{\bm{W}}}\right\|_{F}=\left\|{{\bm{V}}}\right\|_{F}=1}{\mathbb{E}}_{s,r,s^{\prime}}\left[{{\bm{\phi}}(s)^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{W}}{\bm{Y}}(r){\bm{V}}^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s^{\prime})}\right]$
$\qquad\leq\sup_{\left\|{{\bm{W}}}\right\|_{F}=\left\|{{\bm{V}}}\right\|_{F}=1}{\mathbb{E}}_{s,r,s^{\prime}}\left[{\left\|{{\bm{W}}^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s)}\right\|\left\|{{\bm{Y}}(r)}\right\|\left\|{{\bm{V}}^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s^{\prime})}\right\|}\right]$
@@
$\qquad\leq\sqrt{\gamma}\sup_{\left\|{{\bm{W}}}\right\|_{F}=\left\|{{\bm{V}}}\right\|_{F}=1}{\mathbb{E}}_{s,s^{\prime}}\left[{\left\|{{\bm{W}}^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s)}\right\|\left\|{{\bm{V}}^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s^{\prime})}\right\|}\right]$
@@
$\qquad\leq\sqrt{\gamma}\sup_{\left\|{{\bm{W}}}\right\|_{F}=\left\|{{\bm{V}}}\right\|_{F}=1}\sqrt{{\mathbb{E}}_{s}\left[{\left\|{{\bm{W}}^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s)}\right\|^{2}}\right]{\mathbb{E}}_{s^{\prime}}\left[{\left\|{{\bm{V}}^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s^{\prime})}\right\|^{2}}\right]}$
@@
$\qquad=\sqrt{\gamma}\sup_{\left\|{{\bm{W}}}\right\|_{F}=\left\|{{\bm{V}}}\right\|_{F}=1}\sqrt{{\mathbb{E}}_{s}\left[{{\bm{\phi}}(s)^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{W}}{\bm{W}}^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s)}\right]{\mathbb{E}}_{s^{\prime}}\left[{{\bm{\phi}}(s^{\prime})^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{V}}{\bm{V}}^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s^{\prime})}\right]}$
@@
$\qquad=\sqrt{\gamma}\sup_{\left\|{{\bm{W}}}\right\|_{F}=\left\|{{\bm{V}}}\right\|_{F}=1}\sqrt{\operatorname{tr}\left({{\bm{W}}{\bm{W}}^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\mathbb{E}}_{s}\left[{{\bm{\phi}}(s){\bm{\phi}}(s)^{\intercal}}\right]{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}}\right)\operatorname{tr}\left({{\bm{V}}{\bm{V}}^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\mathbb{E}}_{s^{\prime}}\left[{{\bm{\phi}}(s^{\prime}){\bm{\phi}}(s^{\prime})^{\intercal}}\right]{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}}\right)}$
$\qquad=\sqrt{\gamma}\sup_{\left\|{{\bm{W}}}\right\|_{F}=\left\|{{\bm{V}}}\right\|_{F}=1}\sqrt{\operatorname{tr}\left({{\bm{W}}{\bm{W}}^{\intercal}}\right)\operatorname{tr}\left({{\bm{V}}{\bm{V}}^{\intercal}}\right)}$
@@
where we used $\left\|{{\bm{Y}}(r)}\right\|\leq\sqrt{\gamma}$ for any $r\in[0,1]$ by Lemma E.3, and Cauchy-Schwarz inequality.
@@
$\left[{\left({{\bm{C}}^{\intercal}{\bm{C}}}\right)\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}\right]^{-\frac{1}{2}}\bar{{\bm{A}}}^{\intercal}\left[{\left({{\bm{C}}^{\intercal}{\bm{C}}}\right)\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}\right]^{-1}\bar{{\bm{A}}}\left[{\left({{\bm{C}}^{\intercal}{\bm{C}}}\right)\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}\right]^{-\frac{1}{2}}\succcurlyeq\left({1-\sqrt{\gamma}}\right)^{2}.$
D.2 Proof of Lemma 5.2
@@
For simplicity, we omit $t$ in the random variables, for example, we use ${\bm{A}}$ to denote a random matrix with the same distribution as ${\bm{A}}_{t}$. In addition, we omit the subscripts in the expectation, the involving random variables are $s\sim\mu_{\pi},a\sim\pi\left({\cdot\mid s}\right),r\sim{\mathcal{P}}_{R}\left({\cdot\mid s,a}\right),s^{\prime}\sim P\left({\cdot\mid s,a}\right)$.
@@
$\left\|{{\bm{A}}}\right\|\leq$ $\left\|{{\bm{I}}_{K}\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s)^{\intercal}}\right)}\right\|+\left\|{\left({{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right)\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}}\right)}\right\|$
@@
where we used Lemma E.3. Hence, $C_{A}\leq 2(1+\sqrt{\gamma})$.

### v2

@@ Appendix E Omitted Results and Proofs in Section 4
${\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}\preccurlyeq\frac{1}{(1-\sqrt{\gamma})^{2}}\bar{{\bm{A}}}^{\top}\left({{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}^{-1}}\right)\bar{{\bm{A}}}\left({\preccurlyeq\frac{4}{(1-\gamma)^{2}\lambda_{\min}}\bar{{\bm{A}}}^{\top}\bar{{\bm{A}}}}\right),$ (28)
@@
$\left({{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}}\right)\bar{{\bm{A}}}^{\top}\left({{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}^{-1}}\right)\bar{{\bm{A}}}\left({{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}}\right)\succcurlyeq\left({1-\sqrt{\gamma}}\right)^{2}{\bm{I}}_{dK}.$
@@
$\bar{{\bm{A}}}$ $=\left({{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}\right)-{\mathbb{E}}_{s,r,s^{\prime}}\left[{\left({{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right)\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\top}}\right)}\right],$
then for any ${\bm{\theta}}\in{\mathbb{R}}^{dK}$ with $\left\|{{\bm{\theta}}}\right\|=1$,
${\bm{\theta}}^{\top}\left({{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}}\right)\bar{{\bm{A}}}^{\top}\left({{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}^{-1}}\right)\bar{{\bm{A}}}\left({{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}}\right){\bm{\theta}}$
$\qquad=\left\|{\left({{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}}\right)\bar{{\bm{A}}}\left({{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}}\right){\bm{\theta}}}\right\|^{2}$
$\qquad=\left\|{{\bm{\theta}}-{\mathbb{E}}_{s,r,s^{\prime}}\left[{\left({{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right)\otimes\left({{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\top}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}}\right)}\right]{\bm{\theta}}}\right\|^{2}$
$\qquad\geq\left({1-\left\|{{\mathbb{E}}_{s,r,s^{\prime}}\left[{\left({{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right)\otimes\left({{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\top}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}}\right)}\right]{\bm{\theta}}}\right\|}\right)^{2}.$
@@
$\left\|{{\mathbb{E}}_{s,r,s^{\prime}}\left[{\left({{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right)\otimes\left({{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\top}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}}\right)}\right]}\right\|\leq\sqrt{\gamma}.$ (29)
@@
$\left\|{{\mathbb{E}}_{s,r,s^{\prime}}\left[{{\bm{Y}}(r)\otimes\left({{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\top}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}}\right)}\right]}\right\|\leq\sqrt{\gamma}.$
For any vectors ${\bm{w}}=\left({{\bm{w}}(0)^{\top},\cdots,{\bm{w}}(K-1)^{\top}}\right)$ and ${\bm{v}}=\left({{\bm{v}}(0)^{\top},\cdots,{\bm{v}}(K-1)^{\top}}\right)$ in ${\mathbb{R}}^{dK}$, we define the corresponding matrices ${\bm{W}}=\left({{\bm{w}}(0),\cdots,{\bm{w}}(K-1)}\right)$ and ${\bm{V}}=\left({{\bm{v}}(0),\cdots,{\bm{v}}(K-1)}\right)$ in ${\mathbb{R}}^{d\times K}$, then $\left\|{{\bm{w}}}\right\|=\left\|{{\bm{W}}}\right\|_{F}=\sqrt{\operatorname{tr}\left({{\bm{W}}^{\top}{\bm{W}}}\right)}$ and $\left\|{{\bm{v}}}\right\|=\left\|{{\bm{V}}}\right\|_{F}=\sqrt{\operatorname{tr}\left({{\bm{V}}^{\top}{\bm{V}}}\right)}$. With these notations, we have
$\left\|{{\mathbb{E}}_{s,r,s^{\prime}}\left[{{\bm{Y}}(r)\otimes\left({{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\top}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}}\right)}\right]}\right\|$
$\qquad=\sup_{\left\|{{\bm{w}}}\right\|=\left\|{{\bm{v}}}\right\|=1}{\bm{w}}^{\top}{\mathbb{E}}_{s,r,s^{\prime}}\left[{{\bm{Y}}(r)\otimes\left({{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\top}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}}\right)}\right]{\bm{v}}$
$\qquad=\sup_{\left\|{{\bm{w}}}\right\|=\left\|{{\bm{v}}}\right\|=1}{\mathbb{E}}_{s,r,s^{\prime}}\left[{\sum_{i,j=1}^{K}y_{ij}(r){\bm{w}}(i)^{\top}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\top}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{v}}(j)}\right],$
@@
${\bm{w}}(i)^{\top}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\top}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{v}}(j)=\left({{\bm{W}}^{\top}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\top}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{V}}}\right)_{ij},$
@@
$\left\|{{\mathbb{E}}_{s,r,s^{\prime}}\left[{{\bm{Y}}(r)\otimes\left({{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\top}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}}\right)}\right]}\right\|$
$\qquad=\sup_{\left\|{{\bm{W}}}\right\|_{F}=\left\|{{\bm{V}}}\right\|_{F}=1}{\mathbb{E}}_{s,r,s^{\prime}}\left[{\sum_{i,j=1}^{K}y_{ij}(r)\left({{\bm{W}}^{\top}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\top}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{V}}}\right)_{ij}}\right]$
@@
$\qquad=\sup_{\left\|{{\bm{W}}}\right\|_{F}=\left\|{{\bm{V}}}\right\|_{F}=1}{\mathbb{E}}_{s,r,s^{\prime}}\left[{\operatorname{tr}\left({{\bm{Y}}(r){\bm{V}}^{\top}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s^{\prime}){\bm{\phi}}(s)^{\top}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{W}}}\right)}\right]$
$\qquad=\sup_{\left\|{{\bm{W}}}\right\|_{F}=\left\|{{\bm{V}}}\right\|_{F}=1}{\mathbb{E}}_{s,r,s^{\prime}}\left[{{\bm{\phi}}(s)^{\top}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{W}}{\bm{Y}}(r){\bm{V}}^{\top}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s^{\prime})}\right]$
$\qquad\leq\sup_{\left\|{{\bm{W}}}\right\|_{F}=\left\|{{\bm{V}}}\right\|_{F}=1}{\mathbb{E}}_{s,r,s^{\prime}}\left[{\left\|{{\bm{W}}^{\top}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s)}\right\|\left\|{{\bm{Y}}(r)}\right\|\left\|{{\bm{V}}^{\top}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s^{\prime})}\right\|}\right]$
@@
$\qquad\leq\sqrt{\gamma}\sup_{\left\|{{\bm{W}}}\right\|_{F}=\left\|{{\bm{V}}}\right\|_{F}=1}{\mathbb{E}}_{s,s^{\prime}}\left[{\left\|{{\bm{W}}^{\top}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s)}\right\|\left\|{{\bm{V}}^{\top}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s^{\prime})}\right\|}\right]$
@@
$\qquad\leq\sqrt{\gamma}\sup_{\left\|{{\bm{W}}}\right\|_{F}=\left\|{{\bm{V}}}\right\|_{F}=1}\sqrt{{\mathbb{E}}_{s}\left[{\left\|{{\bm{W}}^{\top}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s)}\right\|^{2}}\right]{\mathbb{E}}_{s^{\prime}}\left[{\left\|{{\bm{V}}^{\top}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s^{\prime})}\right\|^{2}}\right]}$
@@
$\qquad=\sqrt{\gamma}\sup_{\left\|{{\bm{W}}}\right\|_{F}=\left\|{{\bm{V}}}\right\|_{F}=1}\sqrt{{\mathbb{E}}_{s}\left[{{\bm{\phi}}(s)^{\top}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{W}}{\bm{W}}^{\top}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s)}\right]{\mathbb{E}}_{s^{\prime}}\left[{{\bm{\phi}}(s^{\prime})^{\top}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{V}}{\bm{V}}^{\top}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s^{\prime})}\right]}$
@@
$\qquad=\sqrt{\gamma}\sup_{\left\|{{\bm{W}}}\right\|_{F}=\left\|{{\bm{V}}}\right\|_{F}=1}\sqrt{\operatorname{tr}\left({{\bm{W}}{\bm{W}}^{\top}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\mathbb{E}}_{s}\left[{{\bm{\phi}}(s){\bm{\phi}}(s)^{\top}}\right]{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}}\right)\operatorname{tr}\left({{\bm{V}}{\bm{V}}^{\top}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\mathbb{E}}_{s^{\prime}}\left[{{\bm{\phi}}(s^{\prime}){\bm{\phi}}(s^{\prime})^{\top}}\right]{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}}\right)}$
$\qquad=\sqrt{\gamma}\sup_{\left\|{{\bm{W}}}\right\|_{F}=\left\|{{\bm{V}}}\right\|_{F}=1}\sqrt{\operatorname{tr}\left({{\bm{W}}{\bm{W}}^{\top}}\right)\operatorname{tr}\left({{\bm{V}}{\bm{V}}^{\top}}\right)}$
@@
where we used $\left\|{{\bm{Y}}(r)}\right\|\leq\sqrt{\gamma}$ for any $r\in[0,1]$ by Lemma G.3, and Cauchy-Schwarz inequality.
@@
$\left({{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}}\right)\bar{{\bm{A}}}^{\top}\left({{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}^{-1}}\right)\bar{{\bm{A}}}\left({{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}}\right)\succcurlyeq\left({1-\sqrt{\gamma}}\right)^{2}{\bm{I}}_{dK}.$
E.2 Proof of Lemma 5.2
@@
For simplicity, we omit $t$ in the random variables, for example, we use ${\bm{A}}$ to denote a random matrix with the same distribution as ${\bm{A}}_{t}$. In addition, we omit the subscripts in the expectation, the involving random variables are $s\sim\mu_{\pi},a\sim\pi\left({\cdot\mid s}\right),(r,s^{\prime})\sim{\mathcal{P}}(\cdot,\cdot\mid s,a)$.
@@
$\left\|{{\bm{A}}}\right\|\leq$ $\left\|{{\bm{I}}_{K}\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s)^{\top}}\right)}\right\|+\left\|{\left({{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right)\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\top}}\right)}\right\|$
@@
where we used Lemma G.3. Hence, $C_{A}\leq 2(1+\sqrt{\gamma})$.